Yield to Maturity problems (is this the correct way to execute these problems?)

a.) If a one-year discount bond that was purchased for $960, pays $1000 at maturity, what is the interest rate if it's held for the entire year?

F-P/P = 1000-960/960= 0.041 = 4.1

b.) If a one-year discount bond that was purchased for $937, pays $1000 at maturity, what is the interest rate if it's held for the entire year?

=6.72
Answer #1

Answer

a)

Formula:

Yield to Maturity for 1 year discount bond = i(interest rate) = (F - P)/P

Here F = 1000 , P = 960

Hence, interest rate = (1000 - 960)/960 = 0.0416 ~ 4.16%

Hence interest rate = 4.16%

b)

Formula:

Yield to Maturity for 1 year discount bond = i(interest rate) = (F - P)/P

Here F = 1000 , P = 937

Hence, interest rate = (1000 - 937)/937 = 0.0672 ~ 6.72%

Hence interest rate = 6.72%